# HiQ-Lip: The First Quantum-Classical Hierarchical Method for Global Lipschitz Constant Estimation of ReLU Networks

### v1

@@ 6.1 Experimental Setup
We conduct experiments on neural networks with varying depths and widths to assess the scalability of HiQ-Lip. The networks are trained using the Adam optimizer for 10 epochs, achieving an accuracy exceeding 93% on the test set. The experiments were conducted on a device with 32GB of memory and an Intel Core i7 12th Gen CPU. The quantum algorithm portion was simulated using Qboson’s Kaiwu SDK, which features 100 qubits. We consider the following network architectures:
@@ 6.2 Comparison Methods
  • GeoLip (Wang et al., 2022): A geometry-based SOTA Lipschitz constant estimation method that provides tight upper bounds.
  • LiPopt (Latorre et al., 2020): An optimization-based SOTA method that computes upper bounds using semidefinite programming.
  • Matrix Product (MP): A naive method that computes the product of the weight matrix norms across layers, providing a loose upper bound.
@@
We focus on estimating the FGL with respect to the output corresponding to the digit 8, as done in previous works (Latorre et al., 2020; Wang et al., 2022). In the result tables, we use “N/A” to indicate that the computation did not finish within a reasonable time frame (over 20 hours). We highlight the time advantage of the quantum approach by bolding the methods with shorter time.
@@ 6.3 Results on Two-Layer Networks
Table 1: Estimated Lipschitz constants for two-layer networks
@@
Table 2: Computation times (in seconds) for two-layer networks
@@
The naive MP method consistently overestimates the Lipschitz constant by three to six times compared to HiQ-Lip, highlighting HiQ-Lip’s advantage in providing tighter upper bounds. The Sampling method yields lower bound estimates below those of HiQ-Lip and GeoLip, confirming that HiQ-Lip effectively captures the upper bound.
@@
Summary: Overall, HiQ-Lip achieves a favorable balance between estimation accuracy and computational efficiency for two-layer networks. It provides tight upper bounds comparable to GeoLip while reducing computation times by up to 2.2x in 256 hidden units network, particularly as the network width increases, demonstrating its scalability and effectiveness. This acceleration is facilitated by leveraging quantum computing capabilities to solve the QUBO formulation efficiently.
@@ 6.4 Results on Multi-Layer Networks
Table 3: Estimated Lipschitz constants for multi-layer networks
Network HiQ-Lip MP A HiQ-Lip MP B GeoLip MP Sampling
Net3 477.47 477.47 465.11 8,036.51 331.12
Net4 3,246.37 1,093.05 923.13 52,420.62 454.73
Net5 26,132.45 1,513.34 1,462.58 327,219.23 547.11
Table 4: Computation times (in seconds) for multi-layer networks
@@
In Table 3, we compare two variants of HiQ-Lip for multi-layer networks: HiQ-Lip MP A and HiQ-Lip MP B. HiQ-Lip MP A uses the coefficient $\frac{1}{2^{d-2}}$ as suggested in previous literature (Latorre et al., 2020), while HiQ-Lip MP B introduces an additional scaling factor, using $\frac{1}{2^{d-2}d^{d-3}}$, to obtain tighter estimates for deeper networks.
@@ 7 Conclusion
We introduced HiQ-Lip, a quantum-classical hybrid hierarchical method for estimating the global Lipschitz constant of neural networks. By formulating the problem as a QUBO and employing graph coarsening and refinement, we effectively utilized CIMs despite current hardware limitations. Our method achieves comparable estimates to existing SOTA methods with faster computation speed up to 2x and 120x in two-layer and multi-layer network. This work highlights the potential of quantum devices in neural network robustness estimation and opens avenues for future research in leveraging quantum computing for complex machine learning problems.
@@ Acknowledgments
This research was supported by the Management Science & Engineering-Bose Quantum Beijing Inc. Foundation 2024.
Appendix A Detailed Algorithm Analysis and Complexity
We provide detailed mathematical derivations of the time complexities for each phase of the HiQ-Lip algorithm.
A.1 Graph Coarsening Phase Complexity
Let $N=n+m$ be the total number of nodes in the initial graph. At each coarsening level $i$, the number of nodes is $N_{i}=\frac{N}{2^{i}}$. The main steps and their complexities are as follows:
  1. 1.
    Node Embedding and Distance Calculation:
    • Embedding all nodes into a $d$-dimensional space: $\mathcal{O}(N_{i}d)$.
    • Computing pairwise distances (optimized using approximate methods): $\mathcal{O}(N_{i}kd)$, where $k$ is a small constant.
  2. 2.
    Node Pair Matching and Merging:
    • Matching nodes (e.g., using greedy algorithms): $\mathcal{O}(N_{i})$.
    • Merging nodes and updating the adjacency matrix: $\mathcal{O}(N_{i}^{2})$.
  3. 3.
    Graph Reduction:
    • Computing the coarser adjacency matrix $A_{i+1}$: $\mathcal{O}(N_{i}^{2})$.
The total time complexity for the coarsening phase is:
$T_{\text{coarsen}}$ $=\sum_{i=0}^{L-1}\mathcal{O}(N_{i}^{2})$
$=\sum_{i=0}^{L-1}\mathcal{O}\left(\left(\frac{N}{2^{i}}\right)^{2}\right)$
$=\mathcal{O}\left(N^{2}\sum_{i=0}^{L-1}\left(\frac{1}{4}\right)^{i}\right)$
$=\mathcal{O}\left(N^{2}\cdot\frac{1-\left(\frac{1}{4}\right)^{L}}{1-\frac{1}{4}}\right)$
$=\mathcal{O}(N^{2})$
Considering $L=\log_{2}N$, the practical time complexity is approximated as $\mathcal{O}(N^{2}\log N)$.
A.2 Graph Refinement Phase Complexity
In the refinement phase, the algorithm improves the solution by mapping it back to finer graphs. At each refinement level $i$, the complexities are:
  1. 1.
    Initialization:
    • Projecting the solution to the finer graph: $\mathcal{O}(N_{i})$.
  2. 2.
    Gain Computation:
    • Computing gains for all nodes: $\mathcal{O}(N_{i}^{2})$.
  3. 3.
    Local Optimization:
    • Solving a QUBO problem of size $K$ using a quantum device: $\mathcal{O}(T_{Q})$, where $T_{Q}$ depends on the quantum hardware.
The total time complexity for the refinement phase is:
$T_{\text{refine}}$ $=\sum_{i=0}^{L-1}\left(\mathcal{O}(N_{i}^{2})+\mathcal{O}(T_{Q})\right)$
$=\mathcal{O}\left(\sum_{i=0}^{L-1}N_{i}^{2}\right)+\mathcal{O}(LT_{Q})$
$=\mathcal{O}(N^{2})+\mathcal{O}(\log N\cdot T_{Q})$
Since $N_{i}=\frac{N}{2^{L-i}}$, we have:
$\sum_{i=0}^{L-1}N_{i}^{2}$ $=\sum_{i=0}^{L-1}\left(\frac{N}{2^{L-i}}\right)^{2}$
$=N^{2}\sum_{k=1}^{L}\left(\frac{1}{2^{k}}\right)^{2}$
$=N^{2}\sum_{k=1}^{L}4^{-k}$
$=N^{2}\cdot\frac{1-4^{-L}}{1-\frac{1}{4}}$
$=\mathcal{O}(N^{2})$
A.3 Quantum Device Complexity
The time $T_{Q}$ required by the quantum device (e.g., CIM) to solve a QUBO problem of size $K$ may be modeled as $T_{Q}=\mathcal{O}(K^{\alpha})$, where $\alpha$ depends on the device’s properties. Since $K$ is much smaller than $N$ and $\alpha$ is a small constant:
$\mathcal{O}(\log N\cdot T_{Q})=\mathcal{O}(\log N\cdot K^{\alpha})$
This term is typically negligible compared to $\mathcal{O}(N^{2}\log N)$. It is widely believed that quantum devices have an exponential speedup compared to classical algorithms, which means that the running time of quantum computers can be ignored in quantum classical hybrid algorithms on large problems.
A.4 Overall Time Complexity
Combining the coarsening and refinement phases:
$T_{\text{total}}$ $=T_{\text{coarsen}}+T_{\text{refine}}$
$=\mathcal{O}(N^{2}\log N)+\mathcal{O}(N^{2})+\mathcal{O}(\log N\cdot T_{Q})$
$=\mathcal{O}(N^{2}\log N)+\mathcal{O}(\log N\cdot T_{Q})$
Assuming $T_{Q}$ is relatively small and $N$ is large, the overall time complexity is dominated by $\mathcal{O}(N^{2}\log N)$.
The HiQ-Lip algorithm’s overall time complexity is:
$T_{\text{total}}=\mathcal{O}(N^{2}\log N)$
where $N=n+m$ is the total number of neurons in the neural network. The algorithm efficiently reduces the problem size through hierarchical coarsening, making it suitable for current quantum hardware limitations.
Appendix B HiQ-Lip for a One-Layer Hidden MLP Network: Theoretical Upper Bound
In this section, we provide a detailed explanation of the theoretical guarantees associated with the HiQ-Lip method for estimating the global Lipschitz constant of a two-layer MLP network. We focus on how HiQ-Lip estimates the $\ell_{\infty}$-FGL, and we describe why the algorithm provides an upper bound for the true Lipschitz constant.
The global Lipschitz constant, $L$, is defined and bounded as shown in Equations (1)-(4) of the main text. For a two-layer fully connected MLP, this is formalized in Equation (5) as:
$L\leq\ell_{\infty}\text{-FGL}=\max_{y\in[0,1]^{n}}\|W^{T}\text{diag}(u)y\|_{q}=\max_{y\in[0,1]^{n}}\|Ay\|_{q},$
where $A=W^{T}\text{diag}(u)$. Utilizing the duality between the $\ell_{1}$ and $\ell_{\infty}$ norms, we transform this continuous formulation into a discrete QUBO problem:
$L\leq\ell_{\infty}\text{-FGL}=H=-\sum_{i=1}^{n}\sum_{j=1}^{m}a_{ij}x_{i}y_{j},$
where the goal is to minimize $H$, which directly corresponds to estimating the global Lipschitz constant. While the graph coarsening and refinement strategies in HiQ-Lip are heuristic, they are designed to maintain the key structural properties of the original graph while ensuring efficient optimization under quantum hardware constraints.
The hierarchical nature of the algorithm introduces a series of approximations at each phase. These approximations result in an upper bound on the optimal solution. Specifically, the solution $H_{\text{HiQ-Lip}}$ obtained by HiQ-Lip satisfies the following relationship:
$\min H\leq H_{\text{HiQ-Lip}},$
indicating that HiQ-Lip provides an upper bound on the true global Lipschitz constant.
B.1 Coarsening Phase
During the coarsening phase, the nodes of the original graph $G_{0}$ are merged to form a coarser graph $G_{c}$, represented by the adjacency matrix $A_{c}$. The edge weights in the coarser graph are aggregated from the finer graph. Let $P$ denote the mapping matrix used to merge the nodes, such that:
$A_{c}=P^{\top}A_{f}P,$
where $A_{f}$ is the adjacency matrix of the finer graph. This aggregation step leads to a loss of detail, meaning not all the edge weights of the finer graph are perfectly preserved in the coarser graph. As a result, the solution $H_{c}$ obtained by solving the QUBO problem on the coarser graph is an approximation of the true minimum $\min H$ on the original graph:
$H_{c}\geq\min H.$
B.2 Quantum Solution on Coarse Graph
When solving the QUBO problem on the coarsest graph $G_{c}$ using a quantum device, the solution obtained, denoted $H_{\text{quantum}}$, is also an approximation due to hardware noise and finite precision. Therefore, we have:
$H_{\text{quantum}}\geq H_{c}.$
B.3 Refinement Phase
In the refinement phase, the solution from the coarsest graph is iteratively mapped back to the finer graphs. Each refinement step involves local optimization, which improves the objective function but does not guarantee that the true optimal solution on the original graph is recovered. Let $H_{f}$ represent the refined solution on the finer graph. Thus, we have:
$H_{f}\geq H_{\text{quantum}}.$
B.4 Final Solution
Finally, the solution $H_{\text{HiQ-Lip}}$ obtained after the refinement phase satisfies the following:
$H_{\text{HiQ-Lip}}\geq H_{f}\geq H_{\text{quantum}}\geq H_{c}\geq\min H.$
This indicates that HiQ-Lip estimates the global Lipschitz constant with an upper bound, and the method is designed to provide a safe approximation, suitable for tasks like robustness certification where an upper bound is preferred.

### v2

@@ 6.1 Experimental Setup
We conduct experiments on neural networks with varying depths and widths to assess the scalability of HiQ-Lip. The networks are trained using the Adam optimizer for 10 epochs, achieving an accuracy exceeding 93% on the test set. All the experiments were conducted with an Xeon(R) Gold 6226R CPU operating at 2.90GHz and 64 GB of RAM. We conducted experiments by using a simulated quantum SDK to solve the QUBO problem and performed additional training on a CIM provided by Beijing QBoson Quantum Technology Co., Ltd. Limited by the difficulty of Qiskit simulations, the results of Qiskit are failed in efficiency.
We consider the following network architectures:
@@ 6.2 Comparison Methods
  • GeoLip (Wang et al. 2022): A geometry-based SOTA Lipschitz constant estimation method that provides tight upper bounds.
  • LiPopt (Latorre et al. 2020): An optimization-based SOTA method that computes upper bounds using semidefinite programming.
  • Weight-Matrix-Norm-Product (MP): A naive method that computes the product of the weight matrix norms across layers, providing a loose upper bound.
@@
We focus on estimating the FGL with respect to the output corresponding to the digit 8, as done in previous works (Latorre et al. 2020; Wang et al. 2022). In the result tables, we use “N/A” to indicate that the computation did not finish within a reasonable time frame (over 20 hours). We highlight the time advantage of the quantum approach by bolding the methods with shorter time.
@@ 6.3 Results on Two-Layer Networks
Table 1: Estimated Lipschitz constants for two-layer networks
@@
Table 2: Computation times (in seconds) for two-layer networks
We have tested on subproblems to estimate the quantum device’s advantage. Using a network with 64 hidden units and subproblem size limited to 500 variables as an example: the simulated annealing solver required an average of 0.5 seconds per solution of the subproblem, while the CIM hardware solver achieved an average solution time of just 0.09 milliseconds. This represents a speed improvement of approximately 5,000 times for CIM hardware, with significantly better QUBO values showing improvements ranging from 8% to 116%. Across twelve test runs on CIM hardware, we consistently obtained superior values compared to quantum simulation SDKs, along with speed improvements exceeding 1,000 times. The CIM solver demonstrated excellent stability: testing the same QUBO matrix five times yielded identical optimal solutions each time. Our simulation experiments have already demonstrated priority, while hardware experiments confirm that noise does not affect performance, with faster computation speeds and tighter numerical results.
@@
The naive MP method consistently overestimates the Lipschitz constant by three to six times compared to HiQ-Lip, highlighting HiQ-Lip’s advantage in providing tighter upper bounds. The Sampling method yields lower-bound estimates below those of HiQ-Lip and GeoLip, confirming that HiQ-Lip effectively captures the upper bound.
@@
Summary: Overall, HiQ-Lip achieves a favorable balance between estimation accuracy and computational efficiency for two-layer networks. It provides tight upper bounds comparable to GeoLip while reducing computation times by up to a 2.2x speedup on networks with 256 hidden units, particularly as the network width increases, demonstrating its scalability and effectiveness. This acceleration is facilitated by leveraging quantum computing capabilities to solve the QUBO formulation efficiently.
@@ 6.4 Results on Multi-Layer Networks
Network HiQ-Lip GeoLip MP Sampling
Net3 0.48 / 0.48 0.47 8.04 0.33
Net4 3.25 / 1.09 0.92 52.42 0.45
Net5 26.13 / 1.51 1.46 327.22 0.55
Table 3: Estimated Lipschitz constants ($\times 10^{3}$) for multi-layer networks. Each HiQ-Lip cell reports A (left) and B (right) settings.
@@
Table 4: Computation times (in seconds) for multi-layer networks
@@
In Table 3, we compare two variants of HiQ-Lip for multi-layer networks: HiQ-Lip MP A and HiQ-Lip MP B. HiQ-Lip MP A uses the coefficient $\frac{1}{2^{d-2}}$ as suggested in previous literature (Latorre et al. 2020), while HiQ-Lip MP B introduces an additional scaling factor, using $\frac{1}{2^{d-2}d^{d-3}}$, to obtain tighter estimates for deeper networks.
@@
7 Related Work
Existing methods for estimating the global Lipschitz constant often rely on relaxations and semidefinite programming (SDP) (Chen et al. 2020; Shi et al. 2022). Although these methods have a solid theoretical foundation, they suffer from high memory consumption and slow computation speeds. Additionally, simpler approaches such as matrix norm approximations, while computationally efficient, tend to provide overly conservative bounds that fail to accurately reflect the network’s true robustness. These limitations highlight the necessity for more efficient and precise methods.
The Formal Global Lipschitz constant (FGL) is a widely studied upper-bound estimation technique, approximating the maximum norm of the gradient operator by assuming all activation patterns are independent and feasible. FGL has been previously adopted in neural network robustness evaluations (Raghunathan et al. 2018; Fazlyab et al. 2019; Latorre et al. 2020; Wang et al. 2022). Nevertheless, direct calculation remains NP-hard, motivating approximation and heuristic methods.
Quantum computing offers tools for NP–hard combinatorial optimization (Khumalo et al. 2022; Choi 2008; Chatterjee et al. 2024; Li et al. 2025b, a). In particular, QUBO–based, quantum-inspired frameworks (e.g., Coherent Ising Machines and quantum annealers) have shown strong performance on graph-structured problems (Inagaki et al. 2016; Date et al. 2019). Graph formulations are broadly effective, e.g., topic-guided graph networks in summarization (Shi and Zhou 2023). However, limited qubits hinder direct quantum deployment on neural networks; decomposition strategies (divide-and-conquer, QUBO simplification to tractable subproblems) have been explored (Pelofske et al. 2021; Zhou et al. 2023), yet their integration into Lipschitz robustness estimation remains under-studied—this work targets precisely that gap.
Our proposed method, HiQ-Lip, explicitly fills this gap by harnessing quantum-classical hybrid optimization. We exploit the equivalence between global Lipschitz constant estimation and the cut-norm problem, enabling the use of quantum optimization methods in neural network robustness evaluation. To our best knowledge, this is the first systematic effort to leverage quantum optimization techniques, specifically CIMs, for global Lipschitz constant estimation in neural networks.
@@ 8 Conclusion
We introduced HiQ-Lip, a quantum-classical hybrid hierarchical method for estimating the global Lipschitz constant of neural networks. By formulating the problem as a QUBO and employing graph coarsening and refinement, we effectively utilized CIMs despite current hardware limitations. Our method achieves comparable estimates to existing SOTA methods with faster computation speed up to 2x and 120x in two-layer and multi-layer networks. This work highlights the potential of quantum devices in neural network robustness estimation and opens avenues for future research in leveraging quantum computing for complex machine learning problems.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62502550, Shenzhen Science and Technology Program (KJZD20240903095700001).